OpenStudy (anonymous):

Given the sequence -1, 2, -4, 8.... What is the 10th term? Write the number only.

OpenStudy (hybrik):

This has a formula to this

OpenStudy (anonymous):

Well first you have to find the common ratio

OpenStudy (jadeishere):

-1 * -2 = 2 2 * -2 = -4 -4 * -2 = 8 8 * -2 = -16 -16 * -2 = 32 you just multiply the last number in the sequence by -2 each time :)

OpenStudy (anonymous):

okay

OpenStudy (hybrik):

-16. 32, -64, 128, -256, 512 is the answer [512]

OpenStudy (anonymous):

There actually is a formula. an = a1*r^(n-1)

OpenStudy (hybrik):

@ComputerNerd Like i said, its a arithmetic sequence

OpenStudy (hybrik):

It has a formula

OpenStudy (hybrik):

a1 is the first number in sequence

OpenStudy (anonymous):

@hybrik its actually a geometric sequence

OpenStudy (anonymous):

Write the common ratio for the sequence: 12,6,3,\frac{3}{2},... Write the number only. I'm stuck on this one as well

OpenStudy (hybrik):

"\frac{3}{2}"?

OpenStudy (anonymous):

3/2

OpenStudy (hybrik):

The ratio is 1/2

OpenStudy (hybrik):

Either you can do this logically or the arithmetic sequence formula

OpenStudy (anonymous):

Identify the 10th term for the sequence: 12,6,3,\frac{3}{2},... Write the answer as an integer or fraction. I'm stuck on this one as well

OpenStudy (anonymous):

it's the same pattern as the other

OpenStudy (hybrik):

@ComputerNerd I noticed when I was using the formula

OpenStudy (anonymous):

XD

OpenStudy (hybrik):

3/4, 3/8, 3/16, 3/32, 3/64, 3/128 is the answer, 3/128, by the way, Geometric sequences have to do with Multiplying/Dividing patterns, Arithmetic has to do with Adding / Subtracting

OpenStudy (jadeishere):

or you could just do it the easy way @Aliypop and do what I said... Lol @ComputerNerd you don't always need to do the formula if there's a simpler way

OpenStudy (anonymous):

@Jadeishere It's always best to know the formula. Your method is okay for terms lower in the sequence but would take too long for terms like 100th term.

OpenStudy (anonymous):

okay